Sir, In triangle ABC if angle B = 90 degrees what is the value Tan A/2 in terms of its sided? Regards, Karthik

Sir,


In triangle ABC if angle B = 90 degrees what is the value Tan A/2 in terms of its sided?


Regards,


Karthik


 

Grade:9

1 Answers

Chetan Mandayam Nayakar
11 years ago

tanA=c/b

let tan(A/2)=x

2x/(1-x2)=c/b,solve for x